# Learning Social Cost Functions for Human-Aware Path Planning

### v1

@@ 5 Experimental Validation
In this section, we first evaluate the generalization of the learned social cost function, and then demonstrate the resulting behaviors in a number of scenarios both in simulation and on a real robot.
@@ 5.1 Learning Evaluation
The network was trained with a generated dataset containing $5\times 10^{5}$ samples for 5 epochs on a machine with an Intel Core i7-9700K CPU and an Nvidia GeForce RTX 2080 Ti, taking approximately 30 minutes. Generated cost maps may differ from the expected test label in a few cells but still produce a socially correct path. For this reason, instead of the accuracy of the cost map on the test set, we measure the success rate of the generated cost maps. We consider a cost map successful if the path induced by the map is correct; that is, it follows the queue appropriately and does not navigate across groups.
@@ 6 Conclusions
This paper proposes a novel method to enhance Social Robotic Navigation by modifying the planner cost function. Differently from past literature, we do not focus on proxemics or estimates of people’s movement based on kinematic information. Instead, we consider areas the robot should plan to avoid based on the position of people in the environment, in accordance to specific social norms. We develop a proof of concept based on two common social tasks: queuing in a line of people and respecting interaction spaces of people talking with each other. We demonstrate the effectiveness of our solution both in simulation and in a real-world scenario, comparing it with traditional goal-based navigation. We focused on learning and navigation, and used off-the-shelf libraries for computer vision. The social maps could be further refined if the vision module was able to detect more than people’s positions, for instance, whether or not they were talking, and with whom.

### v2

@@ 5 Experimental Validation
In this section, we first evaluate the generalization of the learned social cost function, and then demonstrate the resulting behaviors in a number of scenarios both in simulation and on a real robot. Our focus is on evaluating the system’s ability to accurately recognize and handle the predefined static social tasks. Other social factors, such as proxemics, which involve understanding and maintaining appropriate personal space, or navigating dynamic and crowded environments, are not taken into account. Our method is designed as a modular plug-in that can be seamlessly integrated with other planners based on cost maps, which can address any additional aspects of social navigation.
@@ 5.1 Learning Evaluation
The network was trained with a generated dataset containing $5\times 10^{5}$ samples for 5 epochs on a machine with an Intel Core i7-9700K CPU and an Nvidia GeForce RTX 2080 Ti, taking approximately 30 minutes. Although the generated cost maps may show minor discrepancies from the expected correct labels in a few cells, they may still be capable of producing socially acceptable paths. That is, what would increase the error in terms of the accuracy of cost map generation, actually results in a perfectly acceptable behaviour. For this reason, we measure the success rate of the generated cost maps rather than relying solely on the accuracy of the cost map on the test set. A cost map is deemed successful if the path it generates adheres to social norms—specifically, if it follows the queue correctly and avoids navigating through or disrupting groups. Conversely, a cost map is considered unsuccessful if the introduced costs lead to navigation failures or result in paths that violate social norms, such as cutting across groups or deviating from the queue.
@@ 6 Conclusions
This paper proposes a novel method to enhance Social Robotic Navigation by modifying the planner’s cost function. Differently from past literature, we do not focus on proxemics or estimates of people’s movement based on kinematic information. Instead, we consider areas the robot should plan to avoid based on the position of people in the environment, in accordance to specific social norms. We designed a modular plug-in that can add this capability seamlessly to virtually any navigation system. We developed a proof of concept based on two common social tasks: queuing in a line of people and respecting interaction spaces of people talking with each other. We demonstrate the effectiveness of our method both in simulation and in a real-world scenario, comparing it with traditional goal-based navigation. We focus on learning and navigation, and use off-the-shelf libraries for computer vision. The social maps could be further refined if the vision module was able to detect more than people’s positions, for instance, whether or not they were talking, and with whom.
@@
Acknowledgment
The work presented in this paper has born from the collaboration between the PIC4SeR Centre for Service Robotics at Politecnico di Torino and the Sensible Robots Research Group at King’s College London.